# An Atomic Cluster Expansion Potential for Twisted Multilayer Graphene

### v1

@@ III.2.3 Equilibration of twisted trilayer graphene
We perform geometry optimization of trilayer graphene with alternating twist angle of ±$\pm 4.41^{\circ}$. This test is restricted to the ACE (gt) and ACE (lt) potentials. Figure 5 visualizes the two computed equilibrium structures. The potential ACE (lt) trained on locally twisted configurations predicts a realistic equilibrium structure. By contrast, the potential ACE (gt) trained only on globally twisted configurations predicts an equilibrium with noticeable undulations and distortions in the layers, indicating significant interlayer interactions that are incorrectly predicted. The ACE ($\ast$) model shows similar performance to the ACE (lt) model.
@@ III.2.5 Stress-strain curves
We examine several representative systems, including monolayer graphene, bilayer graphene with a twist angle of $4.41^{\circ}$, and trilayer graphene with alternating twist angles of ±$\pm 4.41^{\circ}$. Additionally, we consider bilayer configurations with point defects such as vacancies, Stone–Wales defects, and interlayer interstitials. The resulting stress-strain curves are shown in Figure 7. In the linear elastic regime, the slope of each curve yields the Young’s modulus, $E=\sigma/\epsilon$, characterizing the stiffness of the material. Among the systems studied, the trilayer structure exhibits the highest Young’s modulus, indicating increased stiffness with additional layers. The trends are consistent with DFT reference calculations. We also examine the effect of fine-tuning by incorporating the DFT reference strained configurations into the training set and retraining the model. The resulting fine-tuned model exhibits clear improved agreement with the DFT reference across both test cases.
Figure 7: Stress-strain curves computed using the ACE ($\ast$) potential and the fine-tuned ACE potentials. Left: Monolayer, bilayer (with a twist angle of $4.41^{\circ}$), and trilayer (with alternating twist angles of ±$\pm 4.41^{\circ}$) graphene under uniaxial tension along the armchair direction. Right: Trilayer graphene systems containing different types of point defects, including a single vacancy, a Stone-Wales (SW) defect, and an interlayer interstitial.
@@ III.3 Molecular Dynamics
After systematically examining the accuracy and stability of the fitted ACE models in geometry optimization tasks, we now conduct similar tests of the models’ performance in molecular dynamics (MD) simulations. All molecular dynamics simulations performed in this section were conducted using a Langevin thermostat, targeting the desired temperature with a friction coefficient of 0.01 $\textrm{fs}^{-1}$ and 1fs time-step. As in the previous section, we use twisted bilayer configurations with a twist angle of $4.41^{\circ}$ throughout this section unless otherwise specified.
@@ III.3.1 MD Stability
We perform molecular dynamics simulations of $4.41^{\circ}$ twisted bilayer graphene at 1000K, using the ACE (gt) and ACE (lt) potentials. The results are shown in Figure 8. The MD simulation using ACE (gt) (trained only on globally twisted configurations) becomes unstable after around 2500 time steps. After this point, the temperature rises sharply, accompanied by significant changes in the potential energy and pressure. This confirms our previous observations that the potential energy surface generated by the ACE (gt) model, trained only on globally twisted configurations, contains “unstable regions” that lead to non-physical behavior in the simulations. By contrast, the MD simulation using ACE (lt) (trained on locally twisted configurations) remains stable for the entire trajectory of 20,000 steps, further supporting the value of using locally twisted configurations during model training.
Next, we demonstrate the stability of MD for bilayer graphene in a large domain and much longer simulation time, only for the ACE ($\ast$) potential. Specifically, we analyze thermodynamic stability for a single unit cell of 29.64 twisted bilayer graphene with an interlayer interstitial defect. The total number of atoms in the system is 7397. We perform MD simulation of this system at 1000K for $10^{5}$ time steps. The results are shown in Figure 9. The fast equilibration followed by stable fluctuations of energy, temperature and pressure over this trajectory indicates long-time stability of the MD system.
Figure 9: Left: Atomic configuration of a commensurate unit cell of bilayer graphene twisted by $29.64^{\circ}$, featuring an interlayer interstitial defect, containing a total of 7397 atoms. Right: Evolution of thermodynamic properties at 1000K over $10^{5}$ molecular dynamics steps. The potential exhibits stable molecular dynamics for the entire trajectory.
@@ III.3.2 Dynamical properties
[…]
@@
Figure 10: Left: Diffusion coefficients of monolayer and twisted bilayer graphene for various temperatures. Center: The velocity auto-correlation function (VACF) of twisted bilayer graphene for various temperatures. Right: Thermal conductivity at various temperatures with the fitted ACE ($\ast$) potential for monolayer and twisted bilayer graphene.
The velocity auto-correlation function (VACF) is another important dynamical property, providing insights into the temporal correlations of particle velocities in twisted bilayer graphene. The VACF $C_{v}(t)$ can be computed from the velocities obtained from molecular dynamics simulations and is important for understanding thermal transport and the dynamic stability of the material [75]. The VACF $C_{v}(t)$ is defined as:
@@
In Figure 10, we present the velocity auto-correlation function (VACF) results for twisted bilayer graphene across various temperatures (100K, 300K, 500K, and 700K), plotted on a semi-logarithmic scale. This semi-log plot clearly shows the exponential decay of the VACF over time, with the decay rate increasing as the temperature rises, indicating a faster loss of velocity correlation at higher temperatures. This trend reflects the enhanced atomic vibrations and thermal motion at elevated temperatures, which lead to quicker decorrelation in particle velocities.
@@
In twisted bilayer graphene, these parameters can be significantly altered by the twist angle. Experimental data indicate that the thermal conductivity of TBG is significantly lower than that of monolayer graphene and bilayer graphene. This phenomenon can be attributed to changes in the Brillouin zone and the emergence of additional phonon branches due to the twisted structure, which enhance phonon Umklapp and normal scattering [79, 80].
@@ III.4 Phonon Spectra
Here, $ij$ represent supercell indices for a $5\times 5$ grid of supercells with the central supercell indexed by $i=0$, $\bf R_{j}$ are the supercell lattice vectors, ${\bf r}_{i\mu\alpha}$ the atomic positions, $\bf A$ the atomic structure, and $M$ is the mass of the Carbon atom. The dynamical matrix elements are proportional to the derivatives of the forces. Phonon mode calculations via Density Functional Theory (DFT) typically involve either differentiating the forces by displacing atoms or employing Density Functional Perturbation Theory (DFPT). Both methods are computationally demanding. Both classical and machine learning interatomic potential models can be differentiated analytically to second order at relatively low computational cost.
@@ IV Conclusion
In this study, we have comprehensively explored the use of machine-learning interatomic potentials (MLIPs) to simulate the complex behavior of twisted multilayer graphene, a system characterized by its unique moiré patterns. By leveraging the Atomic Cluster Expansion (ACE) architecture, we developed an interatomic potential model that achieves near-DFT (in-distribution) accuracy while significantly reducing computational costs. Our approach introduces a new method to generate training and testing datasets using locally twisted structures capable of exhaustively sampling the large configuration space of twisted multilayered materials. The model’s performance across a wide range of simulation tasks, including geometry optimizations, molecular dynamics simulations, and phonon calculations, demonstrates the effectiveness of the fitted model both in- and out-of-distribution. Our results demonstrate that the transferrability from non-specialized training datasets to predictions of general twist angles is not guaranteed, but that the locally twisted structures we propose are sufficient to guarantee stable and accurate predictions in a wide range of simulation tasks.
@@ Appendix A Atomic Cluster Expansion (ACE)
We provide a brief review of the Atomic Cluster Expansion (ACE) model but refer to [16, 84, 22, 85, 86, 87] for further details and discussion. Recalling the setup from Section II.2, we consider an atomic configuration $\mathbf{A}:=\{\boldsymbol{r}_{1},\ldots,\boldsymbol{r}_{N}\}\in\mathbb{R}^{3N}$, where $N$ atoms are represented by their position vectors $\boldsymbol{r}_{j}$. The relative position vector between atom $j$ and a reference atom $i$ is $\boldsymbol{r}_{ij}=\boldsymbol{r}_{j}-\boldsymbol{r}_{i}$, and the local atomic environment around atom $i$ with a cutoff radius $r_{\rm cut}$ is $\mathbf{A}_{i}:=\{\boldsymbol{r}_{ij}\}_{j\in\mathcal{N}(i)}$, where $\mathcal{N}(i)$ denotes the set of indices of all atoms within the cutoff radius $r_{\rm cut}$ from atom $i$.
@@
In the above, the $q$-index enumerates the number of all possible invariant couplings through the generalized Clebsch–Gordan coefficients $\mathcal{C}^{{\bm{n}}{\bm{\ell}}q}_{{\bm{m}}}$, :=$\mathcal{M}_{{\bm{\ell}}}:=\{{\bm{m}}\in\mathbb{Z}^{\nu}|-\ell_{\alpha}\leq{\bm{m}}_{\alpha}\leq\ell_{\alpha}\}.$
This parametrization was originally devised in [16]. Its approximation properties and computational complexity are analyzed in detail in [22, 88, 89]. The implementation we employ in our work is described in [64]. The latter reference also described the details of the choice of radial basis $P_{n}$.
@@
:=$\mathcal{B}:=\Big{\{}{\bf B}_{{\bm{n}}{\bm{\ell}}q}\,\Big{|}\,$ $({\bm{n}},{\bm{\ell}})\in\mathbb{N}^{2\nu}\text{ ordered},~{}\sum_{\alpha}{\bm{\ell}}_{\alpha}\text{ even},~{}\sum_{\alpha}{\bm{m}}_{\alpha}=0,$
@@
where the two approximation parameters are the correlation order $\nu_{\rm max}$ and the total degree $D_{\rm tot}$. To facilitate the use of permutation invariance, we define $({\bm{n}},{\bm{\ell}})\in\mathbb{N}^{2\nu}$ as ordered if the vector associated with this tuple is lexicographically ordered [22, Definition 1]. Additionally, we impose the condition that $\sum_{\alpha}{\bm{\ell}}_{\alpha}$ must be even to ensure point reflection symmetry. The constraint $\sum_{\alpha}{\bm{m}}_{\alpha}=0$ arises from the requirement that rotationally invariant basis functions are constructed using specific linear combinations of spherical harmonic products, which couple to a total angular momentum of zero. This property is derived purely within the framework of linear algebra. For a more detailed discussion, we refer the reader to [22, Section 3.4]. With this selection of the basis we can write our ACE parametrization more conveniently as
@@ Appendix B Bayesian Linear Regression
where $\mathbf{O}$ is a vector containing the observation values $E^{\rm ref}$, $F^{\rm ref}$, and $V^{\rm ref}$.$\mathbf{\Psi}$ is the design matrix containing the ACE basis values corresponding to those observations, and $\mathbf{W}$ is a diagonal matrix containing the weights $w_{E,\mathbf{R}}$ and $w_{F,\mathbf{R}}$. $\mathbf{\Gamma}$ specifies the form of the regularizer and $\lambda$ is a scaling parameter determining the relative weight of the regularization. This formulation of the least squares problem is often also called regularized least squares, and the $\lambda\|\mathbf{\Gamma}\mathbf{c}\|^{2}$ term is often called generalized Tikhonov regularization. The default for $\mathbf{\Gamma}$ is zero or the identity, depending on the choice of solver. Our recommendation is to use the smoothness prior introduced in [64] instead for most solvers.
@@
[…]

### v2

@@ III.2.3 Equilibration of twisted trilayer graphene
We perform geometry optimization of trilayer graphene with alternating twist angle of $\pm 4.41^{\circ}$. This test is restricted to the ACE (gt) and ACE (lt) potentials. Figure 6 visualizes the two computed equilibrium structures. The potential ACE (lt) trained on locally twisted configurations predicts a realistic equilibrium structure. By contrast, the potential ACE (gt) trained only on globally twisted configurations predicts an equilibrium with noticeable undulations and distortions in the layers, indicating significant interlayer interactions that are incorrectly predicted. The ACE ($\ast$) model shows similar performance to the ACE (lt) model.
@@ III.2.5 Stress-strain curves
We examine several representative systems, including monolayer graphene, bilayer graphene with a twist angle of $4.41^{\circ}$, and trilayer graphene with alternating twist angles of $\pm 4.41^{\circ}$. Additionally, we consider bilayer configurations with point defects such as vacancies, Stone–Wales defects, and interlayer interstitials. The resulting stress-strain curves are shown in Figure 8. In the linear elastic regime, the slope of each curve yields the Young’s modulus, $E=\sigma/\epsilon$, characterizing the stiffness of the material. Among the systems studied, the trilayer structure exhibits the highest Young’s modulus, indicating increased stiffness with additional layers. The trends are consistent with DFT reference calculations. We also examine the effect of fine-tuning by incorporating the DFT reference strained configurations (60 configurations) into the training set and retraining the model. The resulting fine-tuned model exhibits clear improved agreement with the DFT reference across both test cases.
Figure 8: Stress-strain curves computed using the ACE ($\ast$) potential and the fine-tuned ACE potentials. Left: Monolayer, bilayer (with a twist angle of $4.41^{\circ}$), and trilayer (with alternating twist angles of $\pm 4.41^{\circ}$) graphene under uniaxial tension along the armchair direction. Right: Trilayer graphene systems containing different types of point defects, including a single vacancy, a Stone-Wales (SW) defect, and an interlayer interstitial.
@@ III.3 Molecular Dynamics
After systematically examining the accuracy and stability of the fitted ACE models in geometry optimization tasks, we now conduct similar tests of the models’ performance in molecular dynamics (MD) simulations. All molecular dynamics simulations performed in this section were conducted using a Langevin thermostat, targeting the desired temperature with a friction coefficient of 0.01 $\textrm{fs}^{-1}$ and 1fs time-step. As in the previous section, we use twisted bilayer configurations with a twist angle of $4.41^{\circ}$ with preiodic boundary conditions throughout this section unless otherwise specified. All MD simulations were performed using the Atomic Simulation Environment (ASE) [65].
@@ III.3.1 MD Stability
We perform molecular dynamics simulations of $4.41^{\circ}$ twisted bilayer graphene at 1000K, using the ACE (gt) and ACE (lt) potentials. The results are shown in Figure 14 in Appendix E. The MD simulation using ACE (gt) (trained only on globally twisted configurations) becomes unstable after around 2500 time steps. After this point, the temperature rises sharply, accompanied by significant changes in the potential energy and pressure. This confirms our previous observations that the potential energy surface generated by the ACE (gt) model, trained only on globally twisted configurations, contains “unstable regions” that lead to non-physical behavior in the simulations. By contrast, the MD simulation using ACE (lt) (trained on locally twisted configurations) remains stable for the entire trajectory of 20,000 steps, further supporting the value of using locally twisted configurations during model training.
Next, we demonstrate the stability of MD for bilayer graphene in a large domain and much longer simulation time, only for the ACE ($\ast$) potential. Specifically, we analyze thermodynamic stability for a single unit cell of 29.64 twisted bilayer graphene with an interlayer interstitial defect. The total number of atoms in the system is 7397. We perform MD simulation of this system at 1000K for $10^{5}$ time steps. The results are shown in Figure 9. The fast equilibration followed by stable fluctuations of energy, temperature and pressure over this trajectory indicates long-time stability of the MD system. We further tested the scalability of our MD simulations using a larger commensurate twisted bilayer graphene supercell at $\theta=45.318^{\circ}$ containing 14,884 atoms. The results are showin in Figure 15 in Appendix E.
Figure 9: Left: Atomic configuration of a commensurate unit cell of bilayer graphene twisted by $29.64^{\circ}$, featuring an interlayer interstitial defect, containing a total of 7397 atoms. Right: Evolution of thermodynamic properties at 1000K over $10^{5}$ molecular dynamics steps. The ACE ($\ast$) exhibits stable molecular dynamics for the entire trajectory.
@@ III.3.2 Dynamical properties
where $\bm{r}(t)$ is the particle position at time $t$ and angular brackets $\langle\cdot\rangle$ denotes the canonical average, approximated in practice as an average over all particles in the system and multiple samples of time origins along the MD trajectory.
@@
Figure 10: Left: Diffusion coefficients of monolayer and twisted bilayer graphene for various temperatures. Center: The VACF of twisted bilayer graphene for various temperatures. Right: Thermal conductivity at various temperatures with the fitted ACE ($\ast$) potential for monolayer and twisted bilayer graphene.
The VACF is another important dynamical property, providing insights into the temporal correlations of particle velocities in twisted bilayer graphene. The VACF $C_{v}(t)$ can be computed from the velocities obtained from molecular dynamics simulations and is important for understanding thermal transport and the dynamic stability of the material [70]. The VACF $C_{v}(t)$ is defined as:
@@
In Figure 10, we present the VACF results for twisted bilayer graphene across various temperatures (100K, 300K, 500K, and 700K), plotted on a semi-logarithmic scale. This semi-log plot clearly shows the exponential decay of the VACF over time, with the decay rate increasing as the temperature rises, indicating a faster loss of velocity correlation at higher temperatures. This trend reflects the enhanced atomic vibrations and thermal motion at elevated temperatures, which lead to quicker decorrelation in particle velocities.
@@
In twisted bilayer graphene, these parameters can be significantly altered by the twist angle. The twist introduces a periodic moiré superlattice, which reduces crystal symmetry and enlarges the effective unit cell. This structural modulation folds the phonon bands and leads to the emergence of additional phonon branches. As a result, more phonon-phonon scattering channels become available, especially for Umklapp and normal processes that limit thermal transport. Experimental data indicate that the thermal conductivity of TBG is significantly lower than that of monolayer graphene and bilayer graphene. These enhanced scattering processes reduce the phonon mean free paths and lifetimes, thereby suppressing thermal conductivity. This phenomenon can be attributed to changes in the Brillouin zone and the emergence of additional phonon branches due to the twisted structure, which enhance phonon Umklapp and normal scattering [74, 75].
@@ III.4 Phonon Spectra
Here, $ij$ represent supercell indices for a $5\times 5$ grid of supercells with the central supercell indexed by $i=0$, $\bf R_{j}$ are the supercell lattice vectors, ${\bf r}_{i\mu\alpha}$ the atomic positions, $\bf A$ the atomic structure, and $M$ is the mass of the Carbon atom. The dynamical matrix elements are proportional to the derivatives of the forces. Phonon mode calculations via DFT typically involve either differentiating the forces by displacing atoms or employing Density Functional Perturbation Theory (DFPT). Both methods are computationally demanding. Both classical and machine learning interatomic potential models can be differentiated analytically to second order at relatively low computational cost.
@@ IV Conclusion
In this study, we have comprehensively explored the use of MLIPs to simulate the complex behavior of twisted multilayer graphene, a system characterized by its unique moiré patterns. By leveraging the ACE architecture, we developed an interatomic potential model that achieves near-DFT (in-distribution) accuracy while significantly reducing computational costs. Our approach introduces a new method to generate training and testing datasets using locally twisted structures capable of exhaustively sampling the large configuration space of twisted multilayered materials. The model’s performance across a wide range of simulation tasks, including geometry optimizations, molecular dynamics simulations, and phonon calculations, demonstrates the effectiveness of the fitted model both in- and out-of-distribution. Our results demonstrate that the transferrability from non-specialized training datasets to predictions of general twist angles is not guaranteed, but that the locally twisted structures we propose are sufficient to guarantee stable and accurate predictions in a wide range of simulation tasks.
@@ Appendix A Atomic Cluster Expansion (ACE)
We provide a brief review of the Atomic Cluster Expansion (ACE) model but refer to [17, 23, 59] for further details and discussion. Recalling the setup from Section II.2, we consider an atomic configuration $\mathbf{A}:=\{\boldsymbol{r}_{1},\ldots,\boldsymbol{r}_{N}\}\in\mathbb{R}^{3N}$, where $N$ atoms are represented by their position vectors $\boldsymbol{r}_{j}$. The relative position vector between atom $j$ and a reference atom $i$ is $\boldsymbol{r}_{ij}=\boldsymbol{r}_{j}-\boldsymbol{r}_{i}$, and the local atomic environment around atom $i$ with a cutoff radius $r_{\rm cut}$ is $\mathbf{A}_{i}:=\{\boldsymbol{r}_{ij}\}_{j\in\mathcal{N}(i)}$, where $\mathcal{N}(i)$ denotes the set of indices of all atoms within the cutoff radius $r_{\rm cut}$ from atom $i$.
@@
In the above, the $q$-index enumerates the number of all possible invariant couplings through the generalized Clebsch–Gordan coefficients $\mathcal{C}^{{\bm{n}}{\bm{\ell}}q}_{{\bm{m}}}$, $\mathcal{M}_{{\bm{\ell}}}:=\{{\bm{m}}\in\mathbb{Z}^{\nu}|-\ell_{\alpha}\leq{\bm{m}}_{\alpha}\leq\ell_{\alpha}\}.$
This parametrization was originally devised in [17]. The implementation we employ in our work is described in [59]. The latter reference also described the details of the choice of radial basis $P_{n}$.
@@
$\mathcal{B}:=\Big\{{\bf B}_{{\bm{n}}{\bm{\ell}}q}\,\Big|\,$ $({\bm{n}},{\bm{\ell}})\in\mathbb{N}^{2\nu}\text{ ordered},~\sum_{\alpha}{\bm{\ell}}_{\alpha}\text{ even},~\sum_{\alpha}{\bm{m}}_{\alpha}=0,$
@@
where the two approximation parameters are the correlation order $\nu_{\rm max}$ and the total degree $D_{\rm tot}$. To facilitate the use of permutation invariance, we define $({\bm{n}},{\bm{\ell}})\in\mathbb{N}^{2\nu}$ as ordered if the vector associated with this tuple is lexicographically ordered [23, Definition 1]. Additionally, we impose the condition that $\sum_{\alpha}{\bm{\ell}}_{\alpha}$ must be even to ensure point reflection symmetry. The constraint $\sum_{\alpha}{\bm{m}}_{\alpha}=0$ arises from the requirement that rotationally invariant basis functions are constructed using specific linear combinations of spherical harmonic products, which couple to a total angular momentum of zero. This property is derived purely within the framework of linear algebra. For a more detailed discussion, we refer the reader to [23, Section 3.4]. With this selection of the basis we can write our ACE parametrization more conveniently as
@@ Appendix B Bayesian Linear Regression
where $\mathbf{O}$ is a vector containing the observation values $E^{\rm ref}$, $F^{\rm ref}$, and $V^{\rm ref}$$\mathbf{\Psi}$ is the design matrix containing the ACE basis values corresponding to those observations, and $\mathbf{W}$ is a diagonal matrix containing the weights $w_{E,\mathbf{R}}$ and $w_{F,\mathbf{R}}$. $\mathbf{\Gamma}$ specifies the form of the regularizer and $\lambda$ is a scaling parameter determining the relative weight of the regularization. This formulation of the least squares problem is often also called regularized least squares, and the $\lambda\|\mathbf{\Gamma}\mathbf{c}\|^{2}$ term is often called generalized Tikhonov regularization. The default for $\mathbf{\Gamma}$ is zero or the identity, depending on the choice of solver. Our recommendation is to use the smoothness prior introduced in [59] instead for most solvers.
@@
$p(\mathbf{\mathbf{O}|\Psi},\mathbf{\Sigma}_{0},\beta)$ $=\int p(\mathbf{O}|\mathbf{\Psi},\mathbf{c},\beta)p(\mathbf{c}|\mathbf{\Sigma}_{0})d\mathbf{c}$
@@ Appendix C Training Details
Although this study employs DFT-PBE data, our method is compatible with higher-level reference data (e.g., HSE06, GW), and can be extended accordingly when such data becomes available.
@@ Appendix D Dataset Disregistry Generation
Appendix E Numerical Supplement
We tested the stability of MD simulations at 1000K for $4.41^{\circ}$ twisted bilayer graphene using two ACE models: one trained on globally twisted data (ACE (gt)) and the other on locally twisted data (ACE (lt)). As shown in Figure 14, the ACE (gt) model becomes unstable after 2500 steps, with temperature and energy diverging. In contrast, ACE (lt) remains stable for 20,000 steps, highlighting the importance of locally twisted configurations for robust simulations.
To assess the scalability of our MLIP in large-scale simulations, we performed a molecular dynamics run on a commensurate bilayer graphene system with a twist angle of $45.318^{\circ}$, containing 14,884 atoms and an interlayer interstitial defect. The simulation was carried out at 1000 K using a Langevin thermostat with DFT-D3 dispersion corrections and a time step of 1 fs for 5000 steps. The system remained stable throughout the trajectory, demonstrating the robustness of the ACE ($\ast$)model.
Figure 15: Left: Atomic configuration of a commensurate bilayer graphene structure with a twist angle of $45.318^{\circ}$, containing 14,884 atoms. Right: Evolution of total energy during a 1000 K molecular dynamics simulation over 5000 steps.